A compound containing only C, H, and O, was extracted from the bark of the sassafras tree. The combustion of 23.9 mg produced 64.9 mg of CO2 and 13.3 mg of H2O. The molar mass of the compound was 162 g/mol. Determine its empirical and molecular formulas.
1
Expert's answer
2017-10-06T13:09:07-0400
CxHyOz + O2= xCO2 + 1/2yH2O
n(CO2) = 64.9 / (12 + 16*2) = 1.475 (mol) = n(C)
m(C) = 1.475 * 12 = 17.7 (mg)
n(H2O) = 13.3 / (1*2 + 16) = 0.74 (mol) = 0.5n(H)
m(H) = 0.74 * 2 * 1 = 1.48 (mg)
m(O) = 23.9 – 17.7 – 1.48 = 4.72 (mg)
x/y/z = 17.7/12 / 1.48/1 / 4.72/16 = 1.475 / 1.48 / 0.295 = 5 / 5 / 1 = C5H5O
M(C5H5O) = 12*5 + 5*1 + 16 = 81 (g/mol)
M(CxHyOz)/M(C5H5O) = 162 / 81= 2
C10H10O2
So
Empirical - C5H5O
Molecular - C10H10O2